OpenStudy (anonymous):

How could i make a perimeter of 4 inches for the letter S ?

OpenStudy (john_es):

One solution would be the use of two semicircunferences of equal radius. Then, \[L/2+L/2=\pi R+\pi R=4\Rightarrow R=2/\pi\approx0.64\ \text{ in}\] So you need two semicircunferencies of radius 0.64 in.

OpenStudy (anonymous):

so does that mean in the upper and lower curved part of the S would be .64 in?o:

OpenStudy (john_es):

Exact, each semicircunference will have R=0.64 in, so each semicircunference will have 2 in of perimeter. Total = 4 in of perimeter for letter S.